OpenStudy (anonymous):

Find all solutions in the interval [0, 2π). cos x = sin x x = three pi divided by four., five pi divided by four. x = pi divided by four., seven pi divided by four. x = three pi divided by four, seven pi divided by two x = pi divided by four, five pi divided by four. @sithsandgiggles

Nnesha (nnesha):

welll can you take a screenshot ?

OpenStudy (anonymous):

wont let me

Nnesha (nnesha):

why not ?

Nnesha (nnesha):

check your equation please :-)

OpenStudy (anonymous):

i just can't.. others always help me without a screenshot or anything

Nnesha (nnesha):

alright...

OpenStudy (anonymous):

so can u help me?

OpenStudy (anonymous):

cuz i understand if u cant

Nnesha (nnesha):

yeah i'm going to help you i just wanted to make sure that equation is correct :-)

OpenStudy (anonymous):

alright i can retype it. it's cos x = sin x :-D

Nnesha (nnesha):

alright set it equal to one divide both sides by cos

OpenStudy (anonymous):

i dont know how to do any of this :/

OpenStudy (anonymous):

i'm so sorry

Nnesha (nnesha):

alright there is another easy way to find x values sin(x) = cos (x) in unit circle sin x = y coordinate cos (x) = x coordinate so you can replace sin (x) by y and cos (x) by x x=y now now at what angle x and y values are same ??

Nnesha (nnesha):

http://prntscr.com/7f4tf7

OpenStudy (anonymous):

45 degrees

Nnesha (nnesha):

yes there is one more !

OpenStudy (anonymous):

and 225 degrees

Nnesha (nnesha):

yep you can convert those two angles to radians by using this formula \[\huge\rm d \times \frac{ \pi }{ 180 }\] or you can cheat(use unit circle )

Nnesha (nnesha):

:-)

OpenStudy (anonymous):

lol umm.. x = pi divided by four, five pi divided by four? @Nnesha

Nnesha (nnesha):

yes right

Nnesha (nnesha):

and here is a link https://www.youtube.com/watch?v=ww5pnpVGejk

OpenStudy (anonymous):

WOOOOW! that was easier than i thought!

OpenStudy (anonymous):

thank you so much! may I please get a medal?

Nnesha (nnesha):

np :-)

OpenStudy (anonymous):

:)